Grade 12Physical Chemistry

a microscope using suitable photons is employed to locate an electron in an atom within a distance of 0.1 Å. What is the uncertainty involved in the measurement of its velocity?

sudhanshu
12 Years agoGrade 12

1 Answer

Bharat
8 Years ago
By heisenberg`s Uncertainty Principle ∆x.m.∆V>=h/4π We know: m(mass of electron)=9.11×10^-31kg h =6.626 × 10^-34Js ∆v>=h/4π.m.∆x∆v>= 6.626 ×10^-34 ___________________________________ 4×3.14×9.11×10^-31×0.1×10^-10 ∆v>=0.579×10^7m/s ∆v>=5.79×10^6m/s